### v2

# New binary black hole mergers in the LIGO–Virgo O3b data

Abstract
We report the detection of 6 new candidate binary black hole (BBH) merger signals in the publicly released data from the second half of the third observing run (O3b) of advanced LIGO and advanced Virgo. The LIGO–Virgo–KAGRA (LVK) collaboration reported 35 compact binary coalescences (CBCs) in their analysis of the O3b data Abbott et al. (2021), with 30 BBH mergers having coincidence in the Hanford and Livingston detectors. We confirm 17 of these for a total of 23 detections in our analysis of the Hanford–Livingston coincident O3b data. We identify candidates using a search pipeline employing aligned-spin quadrupole-only waveforms. Our pipeline is similar to the one used in our O3a coincident analysis Olsen et al. (2022), except for a few improvements in the veto procedure and the ranking statistic, and we continue to use an astrophysical probability of one half as our detection threshold, following the approach of the LVK catalogs. Most of the new candidates reported in this work are placed in the upper/lower-mass gap of the black hole (BH) mass distribution. We also identify a possible neutron star-black hole (NSBH) merger. We expect these events to help inform the black hole mass and spin distributions inferred in a full population analysis.
I Introduction
The Advanced LIGO and J Aasi et al. (2015) and Advanced Virgo Acernese et al. (2015) detectors have now been joined by KAGRA Akutsu et al. (2021) to form the LIGO–Virgo–KAGRA (LVK) collaboration, which recently released data from the second half of the third observing run (O3b) to the public Vallisneri et al. (2015). The LVK detected gravitational waves (GW) from 35 compact binary coalescence (CBC) events including two probable neutron star–black hole (NSBH) mergers Abbott et al. (2021). 32 of these were found to be coincident in the Hanford–Livingston (HL) detectors, among which 30 were reported to be binary black hole (BBH) signals. Nitz et al. (2021) independently produced the 4-OGC catalog, which included an additional 7 BBH mergers that were not in the LVK catalog.
In this work, we report the detection of 6 new candidate BBH merger signals from our coincident search of O3b HL data (excluding Virgo). Our search was carried out with a pipeline that is very similar to the one used in Olsen et al. (2022) (see Venumadhav et al. (2019) for an overview of the pipeline). We show these detections in Fig. 1 alongside the 90 CBC signals detected by LVK, additional events detected in 4-OGC, and the events previously detected by our pipeline in O1-O3a with probability of astrophysical origin () greater than one half Abbott et al. (2016a, 2019); Nitz et al. (2019, 2020); Abbott et al. (2021a); The LIGO Scientific Collaboration et al. (2021); Venumadhav et al. (2019, 2020); Olsen et al. (2022).
Understanding the mass and spin distributions of the astrophysical BBH population is one of the major goals of GW astronomy. As the number of events in the catalog increases, statistically inferred measurements of the population parameters become more and more precise Roulet et al. (2021); Abbott et al. (2021b). These improved measurements help us to constrain BBH formation channel rates, and they allow us to probe deeper into the fundamental physics of black holes (BHs) and even cosmology Yang et al. (2017); Perkins et al. (2021); Abbott et al. (2021); O’Brien et al. (2019); Ezquiaga (2021).
One of the standard BBH formation channels consists of binary stars that evolve in isolation. Stellar evolution models predict that, if massive enough, stars can go through pulsational pair-instability (PPISN) and pair-instability supernovae (PISN). In particular, the latter can completely destroy the stars, leaving no remnant at the end of the evolution. The standard prediction is that this will lead to an “upper mass gap” (UMG) in which the astrophysical distribution drops sharply for BH masses between and Fowler and Hoyle (1964); Barkat et al. (1967); Bond et al. (1984); Heger and Woosley (2002); Woosley et al. (2007); Woosley (2017); Farmer et al. (2019); Chen et al. (2014); Yoshida et al. (2016); Belczynski, K. et al. (2016). Local observations of BHs also hint at the existence of an empirical “lower mass gap” (LMG) between and corresponding to the region between the maximum NS mass Alsing et al. (2018) and the lowest observed BH mass Bailyn et al. (1998a); Özel et al. (2010a); Farr et al. (2011a); Belczynski et al. (2012); Farah et al. (2021); Liu et al. (2021) (note that according to Alsing et al. (2018), the maximum mass of the NS can vary between ). Detecting BHs in these gaps can have several implications, e.g., detecting a BH in the UMG suggests an alternative to the isolated stellar evolution mechanism for its formation, such as hierarchical mergers in dense stellar environments (the “dynamical formation” channel) Santoliquido et al. (2020); Di Carlo et al. (2020); Kremer et al. (2020); Mapelli et al. (2021); González et al. (2021); Fabj et al. (2020); McKernan et al. (2020); Tagawa et al. (2021, 2020); Yang et al. (2019); Gröbner, M. et al. (2020)) and/or they can make us revisit the problem of stellar collapse Farmer et al. (2019); Marchant and Moriya (2020); Mehta et al. (2022). As we will see later, among the new events presented here we do find sources whose constituents lie in these gaps (see Table 1). Including such events in the population analysis could impact the inferred BH mass distribution Abbott et al. (2021b); Roulet et al. (2021).
![Refer to caption](x1.png)
![Refer to caption](x2.png)
In addition to the masses, the BH spins can also be measured from GW signals. The best measured parameter, however, is a particular combination of the masses and spins called the effective spin:
where are the masses and are the dimensionless spins projected on the orbital angular momentum. Together with the measurement of the masses, measuring could give us more information about the underlying formation channels. For example, as the total mass of the binaries increases, highly spinning systems with (anti-aligned) become increasingly difficult to form in the isolated binary evolution channel even under a broad range of BH natal kicks Rodriguez et al. (2016). Similar to our previous work Olsen et al. (2022), we employ a spin prior which is uniform in . This prior is also useful if a reader wants to reweight these results to another prior of their choice, e.g., a prior that is isotropic in the individual spins such as the one used in the GWTC-3 Abbott et al. (2021) and 4-OGC Nitz et al. (2021) catalogs . Though we use uniform in as our default spin prior for our parameter estimation runs, we perform an additional set of parameter estimation runs to compare against the isotropic spin prior results.
In this work, we follow-up on our previous study (Olsen et al. (2022)) which looked for new events in the Hanford–Livingston (HL) data from the first half of the third observing run (O3a), and analyze the data from the second half of the third observing run (O3b) to complete an O3 catalog with a similar pipeline (see Section II for a description of the changes). We detect 6 new BBH candidate events under the criterion that the astrophysical probability is at least one half (which is also the threshold used in the GWTC-3 and 4-OGC catalogs). By simply summing the complements of the values assigned to these events, we expect that roughly 2 of the 6 events may be noise transients rather than astrophysical signals (note, however, that the estimates of depend on multiple factors including the astrophysical prior used, and that the dependence is strongest in the case of near-threshold detections such as these).
Our pipeline confirmed the detections of 17 out of the 30 HL coincident BBH mergers reported in the GWTC-3 catalog (thus, in total, we detected 23 candidate CBC events in our coincident search of HL data from O3b). When we investigated our pipeline’s performance on the remaining events in the LVK catalog, we found that 6 of them were marginal in our catalog, i.e., recovered with a lower than one half, while 6 were vetoed for failing one or more of our signal consistency and excess power tests, and one event was missed. Most of the marginally recovered events were detected by only one of the LVK search pipelines, while three of the vetoed events show signs of physical effects like spin-precession and higher harmonics that were not modeled in our template bank. We also note that four of the vetoed events were flagged as having data-quality issues in the GWTC-3 catalog paper. Finally, we also checked that 4 of the 6 new candidates reported in this work were also present in the list of GWTC-3 sub-threshold candidates LIGO Scientific Collaboration and Collaboration (2021).
The new GW sources reported here could come from a range of astrophysical scenarios. Three events have their primary (i.e. more massive) constituents confidently placed in the upper mass gap. One event has its secondary (i.e. less massive) constituent confidently placed in the lower mass gap, i.e. . If we allow for the possibility that the maximum mass of the NS can reach up to Alsing et al. (2018), this event could also be a neutron star-black hole (NSBH) signal at confidence (note however that we did not perform a dedicated NSBH search). There is also a particular low-mass event which, under uniform prior on effective spin, seems to have strongly anti-aligned spins ( at confidence; see the discussion in the text for caveats).
We find that for at least two of the new events, detailed parameter estimation shows some evidence for effects such as higher-order modes (HM) and spin-precession, and also some information from the VIRGO data, which suggests that their significance could be potentially improved by including these elements in future improved searches (our current search of the Hanford–Livingston coincident data uses aligned-spin templates that only have the fundamental () harmonic). As noted earlier, some loud events in the LVK catalog that were vetoed by our pipeline also showed evidence for these physical effects, which additionally motivates the development of improved pipelines incorporating them.
The rest of the paper is organized as follows: in §II we review changes to the IAS pipeline between the O3a and O3b analyses. In §III we discuss the BBH mergers first reported in this work (see Table 1). In §IV we report our results for events already included in GWTC-3 Abbott et al. (2021) and 4-OGC Nitz et al. (2021), noting differences (see Table 2). We summarize the results in §V and discuss the astrophysical implications of the new events. In Appendix A, we show the computation of the astrophysical probability, and also discuss certain details of our veto procedure.
Corner plots of posterior distributions for new events can be found in Appendix B, with PE samples publicly available at https://github.com/AjitMehta/new_BBH_mergers_O3b_IAS_pipeline.
II Changes to the O3a analysis pipeline
Our analysis pipeline is very similar to the version used in our recent analysis of the Hanford–Livingston coincident O3a data Olsen et al. (2022), but we make changes to two components of the pipeline detailed below in this section. Once we have the ranked lists of coincident foreground and background candidates, we calculate the candidates’ values using the same procedure as in Appendix A of Olsen et al. (2022). Note that, unlike the results reported in the previous IAS papers, we also report, for each event, an overall inverse false-alarm rate (IFAR) combined across our template banks (assuming an equal astrophysical rate density for each bank) in addition to the IFAR computed within the respective bank.
II.1 Signal-consistency veto tests:
In the matched-filtering step of our pipeline, we collect and save triggers above a particular threshold on the signal-to-noise ratio (SNR; we use 15 for the template banks for the most massive BBHs, where is the single-detector SNR; the exact threshold varies with the bank and is typically higher for banks for less massive BBHs). In our previous studies, we only performed our signal-consistency veto tests on the subset of these triggers that passed a higher threshold on their SNR (). Ideally, one should perform the veto tests on all the candidates irrespective of their SNR, but we need to pick a higher threshold as the veto tests are computationally expensive (they take roughly a minute per trigger per CPU in a single detector). This choice can lead to some spurious triggers (i.e., ones that should ideally be vetoed) that are below the SNR threshold entering our final ranked lists (this happens to both the coincident foreground and background triggers).
To ameliorate this, we use the following procedure. We first construct our ranked lists using the same veto threshold as before. We then collect high-ranking candidates with SNR below the veto threshold (which had been omitted from our veto tests before), and screen them through the veto tests (we do this procedure on top candidates from our background and coincident lists). We make our final ranked candidate lists by removing any candidates which failed at least one veto test. There is another slight change we make to the order in which the veto procedure is applied to the triggers in the pipeline compared to the O3a pipeline, which we discuss in Appendix A.
II.2 Ranking statistic
Our statistic to rank the triggers performs two functions: it up-weights triggers which have good coherence among multiple detectors, () it down-weights triggers in individual detectors based on the probability with which similar triggers occur in the background (for a detailed discussion on the ranking statistic used in the pipeline, see Appendix A and Ref. Wadekar et al. (2023a)). Corresponding to the second function, our ranking statistic includes a term: – this is the probability of obtaining a trigger with squared SNR from the background for a given detector (in our previous papers, we labelled this term as the rank function (denoted by ), see section J of Ref. Venumadhav et al. (2019)). We measure by constructing a histogram of the SNR values of background triggers (obtained using timeslides) and correcting for the incompleteness introduced by our cuts.
Event | Bank | PE | IFAR (yr) | ||||||||
within bank | overall | ||||||||||
GW200109_195634111This event’s significance could be potentially improved in future searches in which Virgo data is also included in the coincident analysis. | (5, 4) | ||||||||||
GW191228_085854 222This event has bi-modal posterior distributions in the intrinsic parameters. The values quoted here correspond to the first mode. See III.2 for further details. | (1, 0) | ||||||||||
GW200225_075134 | (3, 1) | ||||||||||
GW191117_023843 | (5, 4) | ||||||||||
GW200210_100022 | (5, 2) | ||||||||||
GW200316_235947 | (0, 0) |
In the previous iterations of our search pipeline, we combined background triggers from all templates in a given subbank to construct this histogram and estimate the rank function. In reality, can have different values even for different templates in a given subbank, i.e., the background can be strongly template-dependent, see e.g., Abbott et al. (2016b); Nitz et al. (2017); Wadekar et al. (2023a).
In order to optimize our analysis, we follow the technique in Ref. Wadekar et al. (2023a), where we first cluster the templates within each sub-bank based on their sensitivity to glitches (as quantified empirically by the relative fraction of loud and faint triggers belonging to each template). We then separately construct the histograms of background triggers for each template group, and estimate a rank function for triggers belonging to templates within that group. A detailed discussion on this point will be provided in our subsequent paper Wadekar et al. (2023a).
We expect the senstivity of our pipeline to increase due to these two improvements. We plan to quantify this increase in terms of the volume-time () based on the performance of our pipeline on recovery of injected signals in the data in a future paper.
III Newly Reported BBH mergers
Table 1 lists some properties of the newly reported events: source-frame masses ( and ), effective spin (), redshift (), inverse false alarm rate (IFAR) and estimated .
We perform parameter estimation (PE) using a prior that is uniform in detector-frame constituent masses, effective spin , and comoving volume-time (). The other extrinsic parameters have the standard geometric priors used in GWTC-3 Abbott et al. (2021). We compute redshifts using a cold dark matter (CDM) cosmology with Planck15 parameters Ade et al. (2016). We use the phenomenological waveform approximant IMRPhenomXPHM Pratten et al. (2021), which includes the sub-dominant (or higher-order) modes , and in addition to the dominant mode , and models the effect of orbital precession due to in-plane spins. We accelerate the likelihood evaluations using the relative binning algorithm described in Zackay et al. (2018) as implemented in the PE package Roulet et al. (2022) (see Appendix B for more details). We note that these details are the same as in our O3a search: Olsen et al. (2022).
In the rest of this section, we will briefly comment on the properties of the new events reported in this paper.
III.1 High-mass sources
GW200109_195634
This event is our most significant new detection, with . It has a primary mass () that lies in the UMG (see Fig. 6 below). This event has a redshift () greater than 0.6 at confidence.
We find that the best-fit waveform that the PE explores has a higher value of than the search by , where is the maximum log likelihood achieved within the posterior samples, and represent the SNR recovered by the search template in the Hanford and Livingston detectors. It turns out that a significant fraction of the difference () comes from the Virgo detector which is included in the PE but not in our search. Including Virgo data in the future coincident searches could thus raise the significance of this event.
GW200225_075134
This event, with , has a primary mass measurement of (Fig. 7). It does not pose an issue for the UMG.
GW191117_023843
This event, with , has a primary mass at confidence. Thus, this object could also be a BH in the UMG (Fig. 8).
GW200210_100022
This event is one of the least significant events in our list, with . The PE analyses confidently place the primary component in the UMG, with a measurement of at confidence (Fig. 9). The effective spin (), on the other hand, is measured to be positive with a magnitude greater than at confidence, irrespective of the spin prior used. The mass ratio of the best-fit waveform is also high with .
It might be worth noting that the PE analyses yield a slightly larger than the search; . The additional contribution comes from the HM and precession. This might be expected as the signal has a high mass ratio as well as a high total mass and effective spin. Recall that our search templates are based on the approximant Khan et al. (2016), which only models the dominant mode and does not incorporate precession. Thus, a future search with templates that incorporate these physical effects may further improve the significance of this event.
III.2 Low-mass sources
GW191228_085854
This event, with , has a bi-modal posterior distribution in the masses and effective spin under our default spin prior (see Fig. 12). The first posterior mode occurs at the secondary mass, and effective spin, while the second one occurs at and . Interestingly, the likelihoods of these two modes are very similar, which suggests that the data seems to prefer both modes equally likely. Both these solutions are, however, very interesting astrophysically. While the first solution hints at a highly negatively spinning stellar mass BBH merger, the second solution suggests at a significantly asymmetric () non-spining BBH merger with the secondary mass confidently lying in the LMG.
On the other hand, a PE with LVK prior (i.e., isotropic prior on spins) selects out only the second solution. This might be expected as this prior does not have significant weight at high positive or negative effective spins. In appendix C we discuss this event in further detail and also show similar results (i.e., bi-modal phenomena) from known LVK events in the similar parameter space region when analysed with our default spin prior. This brings out to our attention the usefulness of using a generic prior so that at least the results for other priors of choice can be obtained by properly reweighting the generic result.
![Refer to caption](x3.png)
GW200316_235947
: This event is the most marginal event, with . It has a secondary mass of (Fig. 10), placing its lighter constituent confidently in the LMG. This object is also consistent with being a heavy NS at confidence, if we take the maximum NS mass to be . Based on the likelihood comparison between the regions (NSBH) and (BBH), the NSBH and BBH solutions have comparable support from the data, so the priors may play a deciding role in identifying the secondary object as a BH or NS. The NSBH solution also has a slightly positive effective spin (see Fig. 10), which makes it more interesting for EM followup. In Fig. 2 we show the sky localization of this event. Although the sky localization is not very well constrained, it would be worthwhile to do a follow-up search for an EM counterpart.
IV Comparison to previously reported catalogs
In Table 2, we summarize our pipeline’s results for the O3b Hanford–Livingston coincident events published in the GWTC-3 catalog Abbott et al. (2021). We also include the additional events as well as the significance reported by the 4-OGC catalog Nitz et al. (2021). Our estimation is based on the distribution of foreground and background triggers over the O3b run (see Appendix A). We compute the inverse false-alarm rate independently for each template bank. For a comparison of across different search pipelines, see Section IV of Olsen et al. (2022).
Confidently recovered events
The LVK collaboration has five main search pipelines: cWB Klimenko et al. (2016), MBTA Andres et al. (2022), GstLAL Sachdev et al. (2019), PyCBC and PyCBC_BBH Usman et al. (2016). Whenever the estimated from any one of these pipelines crosses the threshold of one half, the event is declared a confident detection. With this criterion, there were 32 Hanford–Livingston (HL) coincident events in Abbott et al. (2021). However, two of them (GW200115_042309 and GW191219_163120) were NSBH mergers. We did not look for such signals in this work. Of the remaining 30 coincident BBH mergers, our pipeline retained 17 with a significance comparable to or better than the LVK catalog.
Candidates with low
Our pipeline assigned to 6 of the LVK events. Five of these events were detected confidently by only one of the LVK search pipelines: GW191126_115259, GW200208_222617 and GW200308_173609 were detected by only PyCBC_BBH, while GW200306_093714 and GW200322_091133 were detected by only MBTA. Given that each individual pipeline differs in their analysis choices, which can be especially important for near-threshold events like these, it is not surprising that these signals were also considered marginal by our pipeline. The remaining event, GW200202_154313, was detected by more than one pipeline. However, this event also has a substantial SNR contribution from the Virgo detector, which indicates that adding Virgo data to our coincident search may improve its significance in our future analyses.
Vetoed candidates
We implement a series of signal consistency checks in order to reject triggers that are likely to arise from glitches (noise transients) and consequently improve our pipeline’s sensitivity. Noise fluctuations and the incompleteness of our template bank can cause astrophysical signals to be vetoed, but the tests are designed to gain more in true positive rate than the loss they produce from false negatives due to these effects (in a statistical sense).
Six events from the GWTC-3 catalog were vetoed in our search: GW191103_012549, GW191109_010717, GW191113_071753, GW191127_050227, GW200129_065458 and GW200210_092254.
The events GW191109_010717 and GW200129_065458 had fairly high SNR and are high-mass binaries with imprints of precession Abbott et al. (2021). Our search used a template bank that was built using the non-precessing approximant, and the signal consistency tests compare the data only to the best-fit waveform within the template bank – the absence of precession and higher-harmonics in the template can make it more likely for events that exhibit these phenomena to trigger the signal-consistency tests 333We note that some of the vetoed events here are restored in our search with higher-harmonics Wadekar et al. (2023b).. The data for two events was also flagged for data-quality issues in the GWTC3 catalog.
Our pipeline vetoed GW191113_071753 since the residuals after subtracting the best-fit template exhibited excess power beyond the expectation from Gaussian noise in multiple frequency bands. In the GWTC3 analysis, this candidate was marked as affected by transient, non-Gaussian detector noise. The best-fit waveform in PE for this candidate has a significantly asymmetric mass ratio , and the absence of higher modes in our template bank could have also contributed to the failure of the excess power tests.
GW200210_092254 failed “split-tests” designed to check for consistency between the SNR in different sub-ranges of frequency. This event also has a significantly asymmetric mass ratio, but when we investigated further, we found that our pipeline assigned a representative trigger to this event that came from a different region of the template bank than would be implied from the parameters from the inference runs. The template with the correct parameters had a matched-filtering SNR in Hanford that fluctuated below the collection threshold we used to define triggers444We set these thresholds to reduce the large volume of trigger data that we would otherwise need to analyze with the more costly coincidence-section of the pipeline.. In other words, the actual trigger in the Livingston data could not find a coincident one in the Hanford data due to this threshold, and thus the event was assigned to a different bank in which it failed signal consistency tests. We can detect such events by either lowering the SNR threshold for triggering, or performing a targeted search for events with disparate detector-responses, such as in Ref. Zackay et al. (2021).
The other two events were straightforwardly vetoed in our search; GW191127_050227 failed multiple excess power tests in a few specific bands in a unit-second interval, while GW191103_012549 failed the split tests. We also note that the data for GW191127_050227 was flagged as being affected by data-quality issues in the GWTC3 catalog paper.
We further note that, two events GW191129_134029 and GW200202_154313 were recovered with secondary triggers, i.e., their primary triggers were vetoed. This is the reason we recover them with lower SNR, and thus the lower IFAR compared to the LVK analysis (see Table 2).
Lost triggers
Our pipeline missed the event GW200220_061928. This is a high-mass BBH (87 at 6 Gpc) which was only found by PyCBC-Broad, with all other LVK pipelines having no coincident triggers at this time.
Search space excluded from this work
We did not perform a dedicated search for NSBH or BNS signals. We leave this to future work.
V Discussion
Below we present a qualitative discussion of the astrophysical properties of the new events, including where they could possibly contribute to constraints from full population analyses. We plan to conduct a population analysis involving these events in a separate work. We conclude with the summary of our O3b search results and an outline of the major improvements for our upcoming searches unifying O3a and O3b data.
V.1 Astrophysical implications of the new events
The lower mass gap (LMG)
The BH masses inferred from observations of low-mass X-ray binaries Bailyn et al. (1998b); Özel et al. (2010b); Farr et al. (2011b) seem to be concentrated above . Separately, there are observational upper limits on the known NS mass range () Alsing et al. (2018). Thus, there might be no way to form compact objects between and . This is the basis for hypothesizing the LMG in the astrophysical mass distribution. However, it is still unclear whether the existence of such a gap is supported by any solid theoretical or observational evidence (as opposed to, for example, an artifact of systematic effects in the EM observations of x-ray binaries and GW observations of BBH, NSBH, and BNS mergers).
As of now, we have already observed GW sources whose secondary constituents confidently lie below Olsen et al. (2022); Abbott et al. (2021). Population analyses using catalogs that include the LVK O3b events Abbott et al. (2021b); Roulet et al. (2021); Abbott et al. (2023) also do not provide strong evidence for such a gap. Although a substantial reduction in the merger rates above the maximum NS mass () is observed, there is no subsequent rise above (see figure 5 of Abbott et al. (2023)). This indicates that two distinct populations are required for modeling the observed compact objects, but otherwise no such gap is discernible. The detection of more events in this region of parameter space will help us gain a stronger understanding of this question.
Here we present a new event, GW200316_235947, whose secondary constituent might be either a BH in the LMG or a heavy NS. The NS scenario may be more interesting from an astrophysical standpoint. A prior that includes the LMG as a feature would indeed yield only the NSBH solution. This solution would also have a positive effective spin and mass ratio . The positive spin makes it more favorable for an EM counterpart follow-up search, and we show the sky localization for this event in Fig. 2. The BBH solution, , has a slightly negative effective spin with nearly equal mass ratio. Thus, this event looks interesting under either of the potential solutions for the source configuration. Since it has a low , its impact on the population analysis may be limited Roulet et al. (2020). However, the population analysis also recomputes under the various astrophysical priors being tested, so there is a chance that it would become more important in the case of other mass and spin priors.
Another event, GW191228_085854, might have its secondary constituent in the LMG. We saw that, under our default (uniform) prior on the effective spin, this event yielded bi-modal posterior distribution on the secondary mass where the secondary mode lies confidently in the LMG. This solution was also preferred by the isotropic spin prior (LVK prior). This solution also has a high mass ratio, at confidence interval.
The upper mass gap (UMG)
The evolution of massive stars can go through pair-instability phases where the stars are completely disrupted and no remnant is left behind. This produces a gap in the predicted distribution of BH mass. The exact boundary of such a gap is not well determined, as it depends on multiple uncertain factors such as the rate of carbon-to-oxygen burning in the core , angular momentum transport, metallicity, wind loss, etc. Marchant and Moriya (2020); Farmer et al. (2019); Mehta et al. (2022).
Analyses thus far have not found any conclusive evidence for the UMG in the observed BH mass distribution Abbott et al. (2023); Roulet et al. (2021). This either challenges the current theoretical models of stellar evolution, or the gap might be filled by BBH formed via alternative mechanisms such as hierarchical mergers Anagnostou et al. (2020); Fragione et al. (2020); Veske et al. (2021); Fragione et al. (2021). Since the prediction of relative rates of mergers throughout parameter space is highly uncertain, it is hard to tell at this time. As the number of observations in this region of parameter space grows, the structures in the observed BH mass distribution should be able to tell us whether such a gap really exists.
Here we present three new events, GW191117_023843, GW200109_195634 and GW200210_100022, whose primary masses confidently lie between and . Our measurement of the effective spin for the first two events is consistent with zero. However, for the third event (GW200210_100022), the effective spin is measured to be greater than at confidence. The inference holds regardless of the spin prior used, suggesting this event could be a highly positively spinning BH binary. Even for the second event, GW200109_195634, although the Bayesian analysis yields an effective spin measurement consistent with zero, the high likelihood region is concentrated at a positive effective spin (). The spin properties make these events stand out among the existing detections. The significance of these events will likely be improved in future searches where HM and precession are included in the templates and with Virgo data included in the coincident triggering.
Spin
GW signals also allow us to measure the spins of the binary constituents. Specifically, a certain combination of these spins (namely, the effective spin) can be measured more accurately Ajith (2011); Ng et al. (2018). These spin parameters encode crucial information about BBH formation channels. For example, the isolated formation channel (under standard assumptions for parameters like BH natal kicks) is highly unlikely to form anti-aligned or misaligned BBH systems, especially if the total mass is high Rodriguez et al. (2016). The characteristics of BBH spins depend on many physical processes that occur during the evolution of the progenitor stars, e.g., the angular momentum transport Fuller et al. (2019); Fuller and Ma (2019); Belczynski, K. et al. (2020), tides and mass transfer Gerosa et al. (2018); Qin et al. (2019); Bavera, Simone S. et al. (2020), external environments, BH natal kicks Rodriguez et al. (2016), etc. It is expected that BBH formed in isolated environments will have spins preferentially aligned with the orbital angular momentum. On the other hand, BBH formed in dense stellar environments (such as globular clusters) will have their spin directions randomized by interactions, i.e., an isotropic spin distribution Rodriguez et al. (2016); Mandel and O’Shaughnessy (2010).
The GW detectors are slightly biased towards positively aligned () BBH systems due to selection effects.888This is because the systems are relatively louder for fixed mass and distance. This means that we can observe these systems from a larger volume of space. Thus, their fraction in the population of GW observations is expected to be correspondingly larger than their fraction of the true astrophysical population. This is consistent with what we see in Fig. 1. This selection effect is indeed taken into consideration during the population analysis, and thus the inference of astrophysical parameters is not expected to be biased. However, it is not clear that the observation in Fig. 1 is entirely due to selection effects. We know that the anti-aligned systems are shorter in length and hence they could be more prone to being mistaken for glitches. The waveform approximants in this regime might also not be very faithful. One way to address these questions would be to do injection studies where we inject astrophysical signals into the data and let the pipeline discover them. We plan to do this in our next work.
There are some interesting spin properties among the new events reported here which could be informative in the full population analysis. The high-mass binaries GW200210_100022 and GW200109_195634 are likely to have high positive effective spin (). On the other hand, the low-mass binary GW191228_085854 may be a merger with a large negative . If true, this may add valuable information to the full population analysis as there is rarely such an event in this parameter space.
V.2 Concluding remarks
In this work, we report 6 new candidate BBH mergers from the IAS pipeline’s search of O3b Hanford–Livingston coincident data, based on the criteria that the astrophysical probability is greater than one half (following the LVK collaboration Abbott et al. (2021)). We recovered 17 of the 30 Hanford–Livingston coincident BBH events with comparable or improved significance. Our pipeline recovered 6 events with a significance lower than one half; most of these events were detected by only one LVK pipeline. Thus, in total, we present 23 confidently detected events in the Hanford–Livingston coincident data from the second part of LVK’s third observing run (O3b). The new events reported here, except GW200225_075134 and GW191117_023843, are also present in the list of GWTC-3 sub-threshold events released by the LVK collaboration LIGO Scientific Collaboration and Collaboration (2021).
The new events declared here have interesting astrophysical properties. For example, three events have their primary constituents sitting in the UMG, one of which might also be highly positively spinning with at confidence. One event has its secondary constituent placed confidently below the top of the LMG, with a possibility of being an NSBH signal. We also have a low-mass binary which may have strongly anti-aligned spins. As we can see from Fig. 1, at least four events lie in the sparsely populated regions of the – plane. These new events could statistically impact the inferred distributions of BBH component masses and effective spin.
We expect roughly 2 of the newly reported events to be noise transients rather than astrophysical signals, as suggested by summing the complements of the values in Table 1. Nevertheless, we have shown that for at least two events the may be improved in future searches that include Virgo data in coincident analysis, and HM and precession in search templates. The source properties, including significance (), become more sensitive to the prior choices for low-SNR events like ours. To facilitate further analysis and comparisons between priors, we make our posterior samples and other relevant information public at https://github.com/AjitMehta/new_BBH_mergers_O3b_IAS_pipeline. These results can be reweighted with other choices of priors to observe how the results (like the number of events itself) may change. Our prior choices differ from the other catalogs Abbott et al. (2021); Nitz et al. (2021) mainly in the spins. We use a prior that is flat in the effective spin. Our motivation lies in the fact that the effective spin is the best-measured spin parameter, and hence we want to give equal weights to all physically possible values of that parameter Zackay et al. (2019). On the other hand, using an isotropic spin prior will significantly down-weight large effective spins. A full population analysis not only requires the PE samples, but also the of each event Roulet et al. (2020). For this reason, we also provide a file with the full list of triggers, IAS_O3b_triggers.hdf, in the Github repository.
The idea of this work was to complete the search on the full third observing run (O3) data using the same pipeline that was used for our O3a search Olsen et al. (2022). Nonetheless, we already made a few notable improvements in our pipeline for this work, and we continue to plan to make improvements for upcoming searches. In fact, we will soon release results from a search of the full O3 data using templates that include higher harmonics Wadekar et al. (2023b). We also plan to include Virgo detector data in our coincident analysis in future work. Our next immediate task is to do an injection study, which should enable us to assess the sensitivity of our pipeline in different regions of the compact binary parameter space.
Acknowledgements
We thank Horng Sheng Chia, Katerina Chatziioannou for helpful discussions. DW gratefully acknowledges support from the Friends of the Institute for Advanced Study Membership and the Keck foundation. TV acknowledges support from NSF grants 2012086 and 2309360, the Alfred P. Sloan Foundation through grant number FG-2023-20470, the BSF through award number 2022136 and the Hellman Family Faculty Fellowship. BZ is supported by the Israel Science Foundation, NSF-BSF and by a research grant from the Willner Family Leadership Institute for the Weizmann Institute of Science. MZ is supported by NSF 2209991 and NSF-BSF 2207583. This research was also supported in part by the National Science Foundation under Grant No. NSF PHY-1748958. We also thank ICTS-TIFR for their hospitality during the completion of a part of this work.
This research has made use of data, software and/or web tools obtained from the Gravitational Wave Open Science Center (https://www.gw-openscience.org/), a service of LIGO Laboratory, the LIGO Scientific Collaboration and the Virgo Collaboration. LIGO Laboratory and Advanced LIGO are funded by the United States National Science Foundation (NSF) as well as the Science and Technology Facilities Council (STFC) of the United Kingdom, the Max-Planck-Society (MPS), and the State of Niedersachsen/Germany for support of the construction of Advanced LIGO and construction and operation of the GEO600 detector. Additional support for Advanced LIGO was provided by the Australian Research Council. Virgo is funded, through the European Gravitational Observatory (EGO), by the French Centre National de Recherche Scientifique (CNRS), the Italian Istituto Nazionale di Fisica Nucleare (INFN) and the Dutch Nikhef, with contributions by institutions from Belgium, Germany, Greece, Hungary, Ireland, Japan, Monaco, Poland, Portugal, Spain.
Appendix A Vetoes and Estimation of
Let us define “trigger” to mean a template and a segment of data for which the quadrature sum of matched-filter SNR from each detector (let us call it an incoherent score) passes a certain threshold. A typical event will produce triggers above the follow-up threshold associated with multiple templates when the data stream is match-filtered against the template bank (shown in Fig. 3). Within a segment, we pick the trigger with the highest incoherent score as the representative one for further follow-up. At the next stage, the triggers are then assigned a coherent score (see Appendix D of Olsen et al. (2022)) which incorporates information about the consistency of the phases, amplitudes and arrival times between the detectors. These (coherent) scores can be used to rank triggers. However, the different templates within a bank do not represent the same amount of phase-space volume (or prior probability, see figure 18 of Olsen et al. (2022)). Triggers associated with templates from low-probability regions of the bank should be down-weighted relative to those from regions with larger probability in order to improve the search’s overall sensitivity to the population Dent and Veitch (2014). We achieve this by using an additional template prior (see section C of Olsen et al. (2022)). Finally, we combine these two pieces of information to compare the triggers within a few milliseconds of each other across sub-banks or banks. The one which has the highest value of is selected and the event is assigned to that trigger. Different events are then ranked using a statistic which adds the rank function to to account for the non-Gaussian tails from glitches (see subsection III.J of Venumadhav et al. (2019)). More specifically, the final ranking score () used to rank the events is
where the quantity inside the square parenthesis becomes zero for the Gaussian noise distributions. See section 4 of Wadekar et al. (2023a) for further details on the justification of the above equation.
![Refer to caption](extracted/5577231/template_banks_bbh_o3a.png)
Previously, as in Olsen et al. (2022), we applied vetoes on individual triggers before maximizing the statistic over different template banks when picking which bank to assign each trigger to. We found that this procedure led to a slight overproduction of background triggers. Here we give an argument for why that occurs. Say a candidate produced triggers in three subbanks (or banks), and their scores are , and . Suppose further that the trigger associated with was vetoed out when we applied our signal-consistency tests. If were the highest among the three scores, i.e., the “best-fit” template was not a good match to the data, we would not want to keep this candidate in our catalog. However, the maximization process would then pick one of the remaining two lower scores and , and the candidate would enter the catalog. In this work, we switch the order: we first maximize over the scores and then apply vetoes. This way all the three triggers are removed if were the highest and got vetoed. This procedure slightly improves our background. However, we also pay the price of losing a few foreground triggers.
The astrophysical probability () of an event with ranking score is defined as follows;
where denotes the signal hypothesis and the noise hypothesis. Then, and denote the foreground and background trigger densities at the score , respectively. We compute following the same procedure outlined in Appendix B of Olsen et al. (2022). We measure the foreground and background trigger distributions directly from the data. To estimate the background, we generate an additional 2000 O3b runtimes worth of noise realizations from the O3b data using timeslides, i.e., unphysical time shifts between the data at the Hanford and Livingston detectors.
![Refer to caption](x4.png)
Appendix B Posteriors for the new events
Here we present corner plots of the posteriors for a few selected parameters for our new events: the source-frame constituent masses and , the effective spin , and the redshift . The sampling prior used here is the same as in Ref. Olsen et al. (2022), i.e., uniform in detector-frame constituent masses (as in GWTC-3 and 4-OGC), effective spin (for more details on the flat effective spin prior, see Zackay et al. (2019) or Olsen et al. (2021)), and comoving (using a CDM cosmology with Planck15 results Ade et al. (2016)). These samples are publicly available at https://github.com/AjitMehta/new_BBH_mergers_O3b_IAS_pipeline. The posteriors were generated using the same package as in Olsen et al. (2022) but with the most recent commit https://github.com/jroulet/cogwheel/tree/c15d40ab29faf595e6dcd7951b15ccc92cdf1296 at the time when the analysis was conducted. The other details are also the same, e.g., we use 8192 live points with a log-evidence tolerance value of 0.1 for the underlying sampler Buchner, J. et al. (2014).
![Refer to caption](x5.png)
![Refer to caption](x6.png)
![Refer to caption](x7.png)
![Refer to caption](x8.png)
![Refer to caption](x9.png)
![Refer to caption](x10.png)
Appendix C Bi-modality in posteriors: GW191228_085854 and the LVK analog GW200210_092254
One of the new events reported in this work, GW191228_085854, has a posterior distribution that is bi-modal in several intrinsic parameters when we analyse it with our default spin prior that is uniform in the effective spin (See Fig. 12 below, in which this prior is labelled as the “IAS prior”). Posteriors obtained using the LVK collaboration’s default spin prior (isotropic in individual spins), on the other hand, do not show any such feature and pick out only the secondary mode of the result obtained using the IAS prior. The first posterior mode suggests that this system is a BBH merger with large and anti-aligned spins, while the second mode points to a BBH system with significantly asymmetric masses, but zero effective spin.
The fact that the posterior, and consequently the astrophysical interpretation, can differ depending on the prior choices is not new (especially for low SNR events) but this event calls for further investigation. We looked for analogs in the events within the LVK catalog, and found similar (i.e., low total mass) events within. For example, see Fig. 11 for the event GW200210_092254. It can be seen to have very similar features in the posteriors. Other low mass LVK events GW191113_071753 and GW190917_114630 also exhibit bi-modality in some of their intrinsic parameters. We leave detailed investigations of such systems to future work.
![Refer to caption](x11.png)
![Refer to caption](x12.png)
![Refer to caption](x13.png)

### v3

@@
 (2021) independently produced the 4-OGC catalog, which included an additional 7 BBH mergers that were not in the LVK catalog.
-In this work, we report the detection of 6 new candidate BBH merger signals from our coincident search of O3b HL data (excluding Virgo).
-Our search was carried out with a pipeline that is very similar to the one used in Olsen et al.
+However, it also missed 8 events reported in GWTC-3 that have a probability of astrophysical origin () greater than 0.5.
+This emphasizes that different pipelines identify events near the threshold with different significance.
+In this work, we perform a search analysis on the O3b HL coincident data (excluding Virgo) using a pipeline closely resembling the one used in Olsen et al.
 (2022) (see Venumadhav et al.
 (2019) for an overview of the pipeline).
+We report detections of a total of 23 compact binary mergers, 6 of which are reported for the first time here.
+The search did not recover 13 events reported in GWTC-3 and the additional 7 events reported in 4-OGC (more details follow later in the introduction, with a detailed description provided in section IV).
 We show these detections in Fig.
-1 alongside the 90 CBC signals detected by LVK, additional events detected in 4-OGC, and the events previously detected by our pipeline in O1-O3a with probability of astrophysical origin () greater than one half Abbott et al.
+1 alongside the 90 CBC signals detected by LVK, additional events detected in 4-OGC, and the events previously detected by our pipeline in O1-O3a with greater than one half Abbott et al.
 (2016a, 2019); Nitz et al.
@@
 (2021).
-![Refer to caption](x1.png)
-![Refer to caption](x2.png)
 In addition to the masses, the BH spins can also be measured from GW signals. The best measured parameter, however, is a particular combination of the masses and spins called the effective spin:
@@
 By simply summing the complements of the values assigned to these events, we expect that roughly 2 of the 6 events may be noise transients rather than astrophysical signals (note, however, that the estimates of depend on multiple factors including the astrophysical prior used, and that the dependence is strongest in the case of near-threshold detections such as these).
-Our pipeline confirmed the detections of 17 out of the 30 HL coincident BBH mergers reported in the GWTC-3 catalog (thus, in total, we detected 23 candidate CBC events in our coincident search of HL data from O3b).
+Our pipeline confirmed the detections of 17 out of the 30 HL coincident BBH mergers reported in the GWTC-3 catalog.
 When we investigated our pipeline’s performance on the remaining events in the LVK catalog, we found that 6 of them were marginal in our catalog, i.e., recovered with a lower than one half, while 6 were vetoed for failing one or more of our signal consistency and excess power tests, and one event was missed.
@@
 As noted earlier, some loud events in the LVK catalog that were vetoed by our pipeline also showed evidence for these physical effects, which additionally motivates the development of improved pipelines incorporating them.
+We note that the LVK catalog was also produced using aligned-spin templates limited to the harmonic, as in our case.
+However, our veto tests are comparatively more stringent, making it crucial for us to search for signals showing hints of HM or precession using templates that explicitly include these effects.
 The rest of the paper is organized as follows: in §II we review changes to the IAS pipeline between the O3a and O3b analyses.
@@
 GW200109_195634111This event’s significance could be potentially improved in future searches in which Virgo data is also included in the coincident analysis. | (5, 4) | ||||||||||
-GW191228_085854 222This event has bi-modal posterior distributions in the intrinsic parameters.
-The values quoted here correspond to the first mode.
-See III.2 for further details.
-| (1, 0) | ||||||||||
+GW191228_085854222This event has bi-modal posterior distributions in the intrinsic parameters. The values quoted here correspond to the first mode. See III.2 for further details. | (1, 0) | ||||||||||
 GW200225_075134 | (3, 1) | ||||||||||
@@
 The PE analyses confidently place the primary component in the UMG, with a measurement of at confidence (Fig.
-9).
+10).
 The effective spin (), on the other hand, is measured to be positive with a magnitude greater than at confidence, irrespective of the spin prior used.
@@
 This event, with , has a bi-modal posterior distribution in the masses and effective spin under our default spin prior (see Fig.
-12).
+13).
 The first posterior mode occurs at the secondary mass, and effective spin, while the second one occurs at and .
-Interestingly, the likelihoods of these two modes are very similar, which suggests that the data seems to prefer both modes equally likely.
+Interestingly, the likelihoods of these two modes are very similar, which suggests that the data seems to prefer both modes equally.
 Both these solutions are, however, very interesting astrophysically.
@@
 This brings out to our attention the usefulness of using a generic prior so that at least the results for other priors of choice can be obtained by properly reweighting the generic result.
-![Refer to caption](x3.png)
 GW200316_235947
@@
 It has a secondary mass of (Fig.
-10), placing its lighter constituent confidently in the LMG.
+11), placing its lighter constituent confidently in the LMG.
 This object is also consistent with being a heavy NS at confidence, if we take the maximum NS mass to be .
@@
 The NSBH solution also has a slightly positive effective spin (see Fig.
-10), which makes it more interesting for EM followup.
+11), which makes it more interesting for EM followup.
 In Fig.
@@
 (2023b)..
-The data for two events was also flagged for data-quality issues in the GWTC3 catalog.
+The data for two events was also flagged for data-quality issues in the GWTC-3 catalog.
+Currently, the triggers that fail any of our separate veto tests are completely excluded from our catalog.
+In the future, we plan to devise a scheme for including the failed-test candidates back into our catalog, but with an appropriate penalty.
+This approach would ensure that high-significance events are not entirely removed due to the absence of certain physical effects in our templates, though this may come at the cost of slightly increasing the background in our search.
 Our pipeline vetoed GW191113_071753 since the residuals after subtracting the best-fit template exhibited excess power beyond the expectation from Gaussian noise in multiple frequency bands.
-In the GWTC3 analysis, this candidate was marked as affected by transient, non-Gaussian detector noise.
+In the GWTC-3 analysis, this candidate was marked as affected by transient, non-Gaussian detector noise.
 The best-fit waveform in PE for this candidate has a significantly asymmetric mass ratio , and the absence of higher modes in our template bank could have also contributed to the failure of the excess power tests.
@@
 The other two events were straightforwardly vetoed in our search; GW191127_050227 failed multiple excess power tests in a few specific bands in a unit-second interval, while GW191103_012549 failed the split tests.
-We also note that the data for GW191127_050227 was flagged as being affected by data-quality issues in the GWTC3 catalog paper.
+We also note that the data for GW191127_050227 was flagged as being affected by data-quality issues in the GWTC-3 catalog paper.
 We further note that, two events GW191129_134029 and GW200202_154313 were recovered with secondary triggers, i.e., their primary triggers were vetoed.
@@
 The upper mass gap (UMG)
-The evolution of massive stars can go through pair-instability phases where the stars are completely disrupted and no remnant is left behind.
-This produces a gap in the predicted distribution of BH mass.
-The exact boundary of such a gap is not well determined, as it depends on multiple uncertain factors such as the rate of carbon-to-oxygen burning in the core , angular momentum transport, metallicity, wind loss, etc.
+The evolution of massive stars can involve phases such as PPISN and PISN, where the stars either lose a significant fraction of their mass or are completely disrupted, leaving no remnant.
+As a result, PISN create a gap in the BH mass distribution, while PPISN can populate the mass spectrum below this gap.
+The precise boundaries of this gap remain uncertain, as they depend on several factors, including the rate of carbon-to-oxygen burning in the core (), angular momentum transport, metallicity, and wind loss.
 Marchant and Moriya (2020); Farmer et al.
@@
 This selection effect is indeed taken into consideration during the population analysis, and thus the inference of astrophysical parameters is not expected to be biased.
-However, it is not clear that the observation in Fig.
-1 is entirely due to selection effects.
-We know that the anti-aligned systems are shorter in length and hence they could be more prone to being mistaken for glitches.
-The waveform approximants in this regime might also not be very faithful.
-One way to address these questions would be to do injection studies where we inject astrophysical signals into the data and let the pipeline discover them.
-We plan to do this in our next work.
 There are some interesting spin properties among the new events reported here which could be informative in the full population analysis.
@@
 (2023a) for further details on the justification of the above equation.
-![Refer to caption](extracted/5577231/template_banks_bbh_o3a.png)
 Previously, as in Olsen et al.
@@
 We measure the foreground and background trigger distributions directly from the data.
+We show these distributions in Fig.
+5.
 To estimate the background, we generate an additional 2000 O3b runtimes worth of noise realizations from the O3b data using timeslides, i.e., unphysical time shifts between the data at the Hanford and Livingston detectors.
-![Refer to caption](x4.png)
+In Fig. 4 we show as a function of the ranking score . Using the same fitting ansatz as Olsen et al. (2022), we find that
+provides a good fit over the relevant range. Notice that the fit coefficients are different from the ones in Olsen et al. (2022), as we could have expected.
 Appendix B Posteriors for the new events
@@
 (2014).
-![Refer to caption](x5.png)
-![Refer to caption](x6.png)
-![Refer to caption](x7.png)
-![Refer to caption](x8.png)
-![Refer to caption](x9.png)
-![Refer to caption](x10.png)
 Appendix C Bi-modality in posteriors: GW191228_085854 and the LVK analog GW200210_092254
@@
 We leave detailed investigations of such systems to future work.
-![Refer to caption](x11.png)
-![Refer to caption](x12.png)
-![Refer to caption](x13.png)